A cute robot is automatically rolling off the tabletop with an initial speed of 25.0 m/s. In 5 seconds, how far does it travel (a) vertically and (b) horizontally?


1
Expert's answer
2021-03-17T16:52:53-0400

"v_{0x}=25\ (m/s)"

"v_{0y}=0\ (m/s)"


(a)


"y=v_{0y}t+\frac{gt^2}{2}=0\cdot5+\frac{9.8\cdot 5^2}{2}=122.5\ (m)"


(b)


"x=v_{0x}t=25\cdot5=125\ (m)"